OpenStudy (anonymous):

Hi im 8 my mom is sick and I want to know if my ans is right pl. Rachel jogged along a trail that was 1/4 of a mile long. She jogged along the trail 8 times. How many miles did rachel jog in all? I said 4 miles. ?

OpenStudy (anonymous):

Okay. the rail is long 0.25 miles.

OpenStudy (shadowlegendx):

1/4 is 1/4 of 1. So going around the 1/4 trail 4 times will equal 1 If you went around the track 8 times...then how many miles would it have been?

OpenStudy (anonymous):

If Rachel traveled 8 times the rail, what is the distance that she covered?

OpenStudy (anonymous):

Yours will probably be understood better : )

OpenStudy (shadowlegendx):

This can be represented by the equation: \[8\times \left(\begin{matrix}1 \\ 4 \end{matrix}\right) \]

OpenStudy (anonymous):

...

OpenStudy (shadowlegendx):

Going around the track the 1/4 mile long track 8 times

OpenStudy (anonymous):



OpenStudy (shadowlegendx):

o_O

OpenStudy (anonymous):